# Quantum State Preparation with Optimal T-Count

### v1

@@ 1.1 Applications
Finally we remark that it is easy to see that the ${T}$-count upper bound $O(m+\log(1/\varepsilon))$ from Theorem 1.4 is asymptotically tight. On the one hand, the $\Omega(\log(1/\varepsilon))$ ${T}$-count lower bound holds even if $m=1$ [5]. On the other hand, if $U$ is the ${T}$ gate, we cannot hope to compress ${T}^{\otimes m}$ in a generic way, which otherwise contradicts the optimality of Theorem 1.1. See details in Subsection 2.2.
@@ Discussion.
In this work we have characterized the worst-case ${T}$-count of quantum states and diagonal unitaries. A challenging open question is to understand the optimal ${T}$-count for general $n$-qubit unitaries. We can prove an $\widetilde{\Omega}(2^{n})$ ${T}$-count lower bound (see Theorem 4.3 for details). However there is a glaringly large gap between this lower bound and the best known upper bound $\widetilde{O}\left(2^{1.5n}\right)$ [31, 39]. This upper bound is obtained in [31] by reducing the unitary synthesis task to that of preparing $2^{n}$ $n$-qubit states, each of which uses $\widetilde{O}\left(\sqrt{2^{n}}\right)$ ${T}$ gates; [39] also shows how to synthesize a unitary efficiently with $O\left(\sqrt{2^{n}}\right)$ calls to a $2n$-qubit Boolean oracles, each of which needs $O(2^{n})$ ${T}$ gates (see Lemma 2.1).
@@
The remainder of the paper is organized as follows. In Section 2, we establish the optimal ${T}$-count for diagonal unitaries (Theorem 1.2) and in Subsection 2.1 and Subsection 2.2 we describe the applications to batched synthesis (Theorem 1.3) and mass production (Theorem 1.4). In Section 3, we bring in other techniques and prove the optimal ${T}$-count for state preparation (Theorem 1.1). Detailed proofs of lower bounds for Theorem 1.1 (as Theorem 4.1) and Theorem 1.2 (as Theorem 4.2) are presented in Section 4. A canonical form for Clifford+${T}$ circuit with Pauli postselections is proved in Appendix A, which generalizes [5]. An improved analysis of [31] is in Appendix B.
@@ 2 Diagonal unitary synthesis with optimal T-count
In Theorem 1.2, the diagonal entries of the diagonal unitary $D$ can be arbitrary complex numbers of magnitude 1. An important special case is obtained by restricting to “Boolean phase oracles”: diagonal unitaries with ±$\pm 1$ diagonal entries. Such unitaries can be implemented using Clifford+${T}$ circuits with zero error. Moreover, it is known how to achieve this using very few ${T}$ gates [37, 42, 34, 31]. We shall use the following statement which is phrased in terms of the standard oracle that computes the Boolean function into an ancilla register.
@@
Let $b\geq 1$ be an integer and $f\colon\{0,1\}^{n}\to\{0,1\}^{b}$ be an arbitrary Boolean function. Define $U_{f}$ as the unitary mapping $|x\rangle|y\rangle$ to $|x\rangle|y\oplus f(x)\rangle$ for all $x\in\{0,1\}^{n},y\in\{0,1\}^{b}$. Then $U_{f}$ can be implemented exactly by a Clifford+${T}$ circuit using $O\left(\sqrt{b\cdot 2^{n}}\right)$ ${T}$ gates and ancillary qubits.
@@
We sketch the idea to prove Lemma 2.1. Assume $f$ has a classical circuit consisting of XOR gates and AND gates with bounded fan-in. By converting XOR gate into CNOT gate (which is Clifford) and AND gate into Toffoli gate (which has a constant ${T}$-count), we can synthesize $U_{f}$ with ${T}$-count proportional to the number of AND gates. The latter quantity is termed “multiplicative complexity” in the classical complexity theoretic literature and is thoroughly studied in e.g., [37, 42]; here we also give a brief overview of the upper bound.
Assume $b=1$ for simplicity. One way to compute $f$ is to gradually fix each input bit: $f(x)=\left(x_{1}\land f_{1}(x_{2},\ldots,x_{n})\right)\oplus f_{2}(x_{2},\ldots,x_{n})$, where $f_{2}$ (resp., $f_{1}\oplus f_{2}$) is $f$ with $x_{1}$ fixed to 00 (resp., $1$). Let $d\in[n]$ be an integer to be optimized later. If we iterate the above expansion for $x_{1},\ldots,x_{d}$, then we need $2^{d}$ AND gates and it remains to compute $2^{d}$ Boolean functions on $x_{d+1},\ldots,x_{n}$. Now observe that once all conjunctions $\bigwedge_{j\in S}x_{j}$, for $S\subseteq\left\{d+1,\ldots,n\right\}$ are computed (which takes $2^{n-d}$ AND gates [^8] To see this, note that each $|S|=k$ is an AND of some size-$(k-1)$ conjunction and one extra variable. Therefore, building on size-$(<k)$ conjunctions, we only need to introduce $\binom{n-d}{k}$ AND gates to get size-$(\leq k)$ conjunctions. So the total cost is $\sum_{k\geq 2}\binom{n-d}{k}<2^{n-d}$.), we just need to XOR some of them for each one of the remaining $2^{d}$ Boolean functions. Therefore, the total number of AND gates is $2^{d}+2^{n-d}$, which has the claimed minimum value $O(\sqrt{2^{n}})$ by setting $d=n/2$.
We shall also use the known ${T}$-optimal synthesis of single-qubit unitaries without ancillas [27, 43, 28, 41].
@@
Any single-qubit unitary with determinant 1111999As noted in [43], this determinant condition is needed since a Clifford+${T}$ circuit has determinant in $\{e^{i\cdot t\pi/4}\colon t=0,1,\ldots,7\}$. This is not a problem in practice because the determinant can be adjusted by multiplying the unitary with a global phase. can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using $O(\log(1/\varepsilon))$ ${T}$ gates and without ancillary qubits.
Below we give the proof of Theorem 1.2. The idea is to view the diagonal unitary $D$ as a single-qubit unitary acting on the $n$th qubit, controlled on the state of the first $n-1$ qubits. For each fixed value of the control bits, we have a single-qubit unitary that can be approximated by a sequence [^10] Since Lemma 2.3 uses no ancilla, the Clifford gates are only Hadamard ${H}$ and Phase gate $S=T^{2}$. Hence we simply assume the sequence contains only Hadamard and ${T}$ gates. of Hadamard and ${T}$ gates using Lemma 2.3. To implement $D$, we first use Lemma 2.1 to compute a Boolean function that takes as input the $n-1$ control bits, and outputs a description of the approximating sequence of Hadamard and ${T}$ gates into an ancilla register. Controlled on this ancilla register, we then apply the corresponding sequence of single-qubit gates.
@@
For each $j=1,2,\ldots,2^{n}$, define $U_{j}=\mathsf{diag}(\alpha_{j},\overline{\alpha}_{j})$, which is a single-qubit unitary with determinant $1$. Then $D^{\prime}=\mathsf{diag}(U_{1},\ldots,U_{2^{n}})$. Moreover, we know from Lemma 2.3 that each $U_{j}$ is $\varepsilon$-close to a single-qubit unitary $\widetilde{U}_{j}$ that can be exactly implemented as a sequence of $O(\log(1/\varepsilon))$ Hadamard and ${T}$ gates. It follows that $\|U-\widetilde{U}\|\leq\varepsilon$, where $\widetilde{U}=\mathsf{diag}(\widetilde{U}_{1},\ldots,\widetilde{U}_{2^{n}})$. To complete the proof we describe a Clifford+${T}$ circuit that exactly implements $\widetilde{U}$ using the number of ${T}$ gates and ancillas claimed in the Theorem statement.
@@
Then by Lemma 2.1, the corresponding $(n+2K)$-qubit unitary $F$ mapping $|j\rangle|y\rangle$ to $|j\rangle|y\oplus f(j)\rangle$ can be implemented with $O\left(\sqrt{K\cdot 2^{n}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ ${T}$ gates and ancillas.
Once the gate sequence is computed into an ancilla register, it remains to apply ${H}$ or ${T}$ on the $(n+1)$-th qubit controlled on this ancilla register. For this purpose, we note that controlled-${H}$ and controlled-${T}$ are two-qubit unitaries that can be exactly implemented with $O(1)$ ${T}$ gates without ancillas [20].
@@
    For each $\ell=1,2,\ldots,2K$, we either apply ${H}$ on $\mathcal{A}_{1}$ controlled by $\mathcal{B}_{\ell}$ (if $\ell$ is odd), or we apply ${T}$ on $\mathcal{A}_{1}$ controlled by $\mathcal{B}_{\ell}$ (if $\ell$ is even).
@@
The number of ${T}$ gates used in the above implementation is upper bounded as
@@ 2.1 Batched synthesis
First we discuss batched synthesis of single-qubit unitaries: we will see that we can implement $m=O(\log\log(1/\varepsilon))$ single-qubit gates to error $\varepsilon$ with ${T}$-count $O(\log(1/\varepsilon))$.
We are interested in synthesizing a general tensor product $U_{1}\otimes U_{2}\otimes\cdots\otimes U_{m}$ of single-qubit unitaries. In order to use Theorem 1.2 we will first reduce to the case where each $U_{j}$ is diagonal in the standard basis. To this end we can use the well-known Euler angle decomposition (see e.g., [36]) which expresses a single-qubit unitary $U$ as a product $U=e^{i\delta}e^{i\gamma{Z}}e^{i\beta{X}}e^{i\alpha{Z}}$ for some real numbers $\alpha,\beta,\gamma,\delta$. Here ${X},{Y},{Z}$ are the Pauli matrices. Since $e^{i\beta{X}}={H}e^{i\beta{Z}}{H}$ we can summarize this fact as follows.
@@
Let $U$ be a single-qubit unitary. Then there exist single-qubit diagonal unitaries $A,B,C$ such that $U=A{H}B{H}C$, where ${H}$ is the single-qubit Hadamard matrix.
@@
Let $m\geq 1$ be an integer and $U_{1},U_{2},\ldots,U_{m}$ be arbitrary single-qubit unitaries. Define an $m$-qubit unitary $U=U_{1}\otimes\cdots\otimes U_{m}$. Then $U$ can be implemented up to error $\varepsilon$ by a Clifford+${T}$ circuit using $O\left(\sqrt{2^{m}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$ ${T}$ gates and ancilla qubits.
@@
By Fact 2.4, each $U_{j}$ can be expanded as $A_{j}{H}B_{j}{H}C_{j}$ where $A_{j},B_{j},C_{j}$ are single-qubit diagonal unitaries. As a result, we have $U=A{H}^{\otimes m}B{H}^{\otimes m}C$, where $A=A_{1}\otimes\cdots\otimes A_{m},B=B_{1}\otimes\cdots\otimes B_{m},C=C_{1}\otimes\cdots\otimes C_{m}$ are $m$-qubit diagonal unitaries, which can each be implemented using Theorem 1.2 with the claimed number of ${T}$ gates and ancillas. ∎
@@ 2.2 Mass production
Before proving this, let us describe our high-level strategy. We can use Fact 2.4 to reduce to the special case of implementing a diagonal single-qubit unitary $A:=\mathsf{diag}(1,e^{i\theta})$, by choosing the (irrelevant) global phase to be such that all 3 diagonal unitaries are of this form. For any computational basis input $|x\rangle$ where $x\in\{0,1\}^{m}$, we have $A^{\otimes m}|x\rangle=e^{i\theta\cdot|x|}|x\rangle$, where $|x|$ is the Hamming weight of $x$. A natural approach is then to first compute the Hamming weight into a second register containing $\lceil\log m\rceil$ qubits, then apply the corresponding phase controlled on this register. The Hamming weight can be computed using $O(m)$ gates, as we discuss below. In the second step we need to implement a diagonal unitary on $\lceil\log m\rceil$ qubits. We can do this using Theorem 1.2, incurring a ${T}$-count $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)=O\left(m+\log(1/\varepsilon)\right)$.
@@
where $|x|\in\left\{0,1,\ldots,m-1\right\}$ is the Hamming weight of $x$. Then $U_{\mathsf{Ham}}$ can be exactly prepared by Clifford+${T}$ circuit using $O(m)$ total gates and ancillary qubits.
We also remark that [11] gives the exact multiplicative complexity of computing the Hamming weight function, which upper bounds the ${T}$-count of implementing $U_{\sf Ham}$ (using the connection in Remark 2.2). Their construction has slightly lower T-count than Fact 2.6.
@@
We describe a classical Boolean circuit with $O(m)$ gates to compute the Hamming weight of an $m$-bit string. Here the circuit contains AND, OR, NOT gates with bounded fan-in. Then we can convert this to a reversible classical circuit in the standard way using Toffoli gates and $O(m)$ ancilla qubits; implementing each Toffoli using $O(1)$ ${T}$ gates gives Fact 2.6.
Firstly note that given two $t$-bit numbers, we can compute their sum as a $(t+1)$-bit number using $O(t)$ gates, by the textbook algorithm sequentially computing each bit and the corresponding carry bit. The Boolean circuit to compute the Hamming weight will use the above observation iteratively for $t=1,2,\ldots,\log m$ as follows. We view the input as $m$ $1$-bit numbers. Then we partition it into $m/2$ pairs and compute the sum of each pair as a $2$-bit number. Continuing in this way, at the $t$-th stage we will have $m/2^{t}$ $t$-bit numbers. The total gate complexity is
@@
By Fact 2.4, we obtain single-qubit diagonal unitaries $A,B,C$ such that $U=A{H}B{H}C$. Thus $U^{\otimes m}=A^{\otimes m}{H}^{\otimes m}B^{\otimes m}{H}^{\otimes m}C^{\otimes m}$. Therefore it suffices to show how to implement $A^{\otimes m}$.
@@
    This uses $O(m)$ ${T}$ gates and ancillary qubits by Fact 2.6.
@@
    This uses $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\leq O(m+\log(1/\varepsilon))$ ${T}$ gates and ancilla qubits by Theorem 1.2.
@@
    This again uses $O(m)$ ${T}$ gates and ancilla qubits by Fact 2.6.
Finally we prove the tightness of Theorem 1.4. On the one hand, the $\Omega(\log(1/\varepsilon))$ ${T}$-count lower bound holds even if $m=1$ [5]. On the other hand, consider the special case where $U$ is the ${T}$ gate. Then $U^{\otimes m}={T}^{\otimes m}$ and it is impossible to implement it using $o(m)$ ${T}$ gates. Otherwise, by gate injection $|{T}\rangle^{\otimes m}$ can be approximated with ${T}$-count $o(m)$. Plugging this in Theorem 1.1, we have that any $n$-qubit state can be approximated up to any constant error with ${T}$-count $o(\sqrt{2^{n}})$, contradicting the lower bound for Theorem 1.1.111111The lower bound for Theorem 1.1 holds in the presence of Pauli postselections, as needed here after gate injection. See details in Section 4.

### v2

@@ 1.1 Applications
Finally we remark that it is easy to see that the T{T}-count upper bound $O(m+\log(1/\varepsilon))$ from Theorem 1.4 is asymptotically tight. On the one hand, the $\Omega(\log(1/\varepsilon))$ T{T}-count lower bound holds even if $m=1$ [5]. On the other hand, if $U$ is the T{T} gate, we cannot hope to compress ${T}^{\otimes m}$ in a generic way, which otherwise contradicts the optimality of Theorem 1.1. See details in Subsection 2.2.
@@ Discussion.
In this work we have characterized the worst-case T{T}-count of quantum states and diagonal unitaries. A challenging open question is to understand the optimal T{T}-count for general $n$-qubit unitaries. We can prove an $\widetilde{\Omega}(2^{n})$ T{T}-count lower bound (see Theorem 4.3 for details). However there is a glaringly large gap between this lower bound and the best known upper bound $\widetilde{O}\left(2^{1.5n}\right)$ [31, 39]. This upper bound is obtained in [31] by reducing the unitary synthesis task to that of preparing $2^{n}$ $n$-qubit states, each of which uses $\widetilde{O}\left(\sqrt{2^{n}}\right)$ T{T} gates; [39] also shows how to synthesize a unitary efficiently with $O\left(\sqrt{2^{n}}\right)$ calls to a $2n$-qubit Boolean oracles, each of which needs $O(2^{n})$ T{T} gates (see Lemma 2.1).
While our focus is on $T$-count and our construction gives the asymptotic optimal bound, the number of Clifford gates in our state preparation circuit is $O(2^{n}\log(1/\varepsilon))$, which can be verified given the explicit description of our synthesis procedure. This bound is close to the optimal $\Omega(2^{n}/n)$ lower bound. In particular, if the goal is to minimize total gate count, we can replace Lemma 2.1 and Remark 2.2 with the construction that achieves the $O(2^{n}/n)$ gate count for Boolean functions; then our final circuit will have $O(2^{n}/n)$ gates, albeit a similar amount of $T$ gates. It remains an open problem whether it is possible to achieve $O(\sqrt{2^{n}})$ $T$-count and $O(2^{n}/n)$ total gate count at the same time.
Subsequent Work.
A recent followup by Tan [44] showed how to improve the unitary synthesis $T$-count to $\widetilde{O}(2^{4n/3})$. This is a significant improvement over the aforementioned $\widetilde{O}(2^{1.5n})$ bound, but is still far from the $\widetilde{\Omega}(2^{n})$ lower bound.
@@
The remainder of the paper is organized as follows. In Section 2, we establish the optimal T{T}-count for diagonal unitaries (Theorem 1.2) and in Subsection 2.1 and Subsection 2.2 we describe the applications to batched synthesis (Theorem 1.3) and mass production (Theorem 1.4). In Section 3, we bring in other techniques and prove the optimal T{T}-count for state preparation (Theorem 1.1). Detailed proofs of lower bounds for Theorem 1.1 (as Theorem 4.1) and Theorem 1.2 (as Theorem 4.2) are presented in Section 4. A canonical form for Clifford+T{T} circuit with Pauli postselections is proved in Appendix A, which generalizes [5]. An improved analysis of [31] is in Appendix B.
@@ 2 Diagonal unitary synthesis with optimal T-count
In Theorem 1.2, the diagonal entries of the diagonal unitary $D$ can be arbitrary complex numbers of magnitude 1. An important special case is obtained by restricting to “Boolean phase oracles”: diagonal unitaries with $\pm 1$ diagonal entries. Such unitaries can be implemented using Clifford+T{T} circuits with zero error. Moreover, it is known how to achieve this using very few T{T} gates [37, 42, 34, 31]. We shall use the following statement which is phrased in terms of the standard oracle that computes the Boolean function into an ancilla register.
@@
Let $b\geq 1$ be an integer and $f\colon\{0,1\}^{n}\to\{0,1\}^{b}$ be an arbitrary Boolean function. Define $U_{f}$ as the unitary mapping $|x\rangle|y\rangle$ to $|x\rangle|y\oplus f(x)\rangle$ for all $x\in\{0,1\}^{n},y\in\{0,1\}^{b}$. Then $U_{f}$ can be implemented exactly by a Clifford+T{T} circuit using $O\left(\sqrt{b\cdot 2^{n}}\right)$ T{T} gates and ancillary qubits. [^12] This construction uses $O(b\cdot 2^{n})$ Clifford gates.
@@
We sketch the idea to prove Lemma 2.1. Assume $f$ has a classical circuit consisting of XOR gates and AND gates with bounded fan-in. By converting XOR gate into CNOT gate (which is Clifford) and AND gate into Toffoli gate (which has a constant T{T}-count), we can synthesize $U_{f}$ with T{T}-count proportional to the number of AND gates. The latter quantity is termed “multiplicative complexity” in the classical complexity theoretic literature and is thoroughly studied in e.g., [37, 42]; here we also give a brief overview of the upper bound.
Assume $b=1$ for simplicity. One way to compute $f$ is to gradually fix each input bit: $f(x)=\left(x_{1}\land f_{1}(x_{2},\ldots,x_{n})\right)\oplus f_{2}(x_{2},\ldots,x_{n})$, where $f_{2}$ (resp., $f_{1}\oplus f_{2}$) is $f$ with $x_{1}$ fixed to 0 (resp., $1$). Let $d\in[n]$ be an integer to be optimized later. If we iterate the above expansion for $x_{1},\ldots,x_{d}$, then we need $2^{d}$ AND gates and it remains to compute $2^{d}$ Boolean functions on $x_{d+1},\ldots,x_{n}$. Now observe that once all conjunctions $\bigwedge_{j\in S}x_{j}$, for $S\subseteq\left\{d+1,\ldots,n\right\}$ are computed (which takes $2^{n-d}$ AND gates [^13] To see this, note that each $|S|=k$ is an AND of some size-(k1)(k-1) conjunction and one extra variable. Therefore, building on size-$(<k)$ conjunctions, we only need to introduce $\binom{n-d}{k}$ AND gates to get size-$(\leq k)$ conjunctions. So the total cost is $\sum_{k\geq 2}\binom{n-d}{k}<2^{n-d}$.), we just need to XOR some of them for each one of the remaining $2^{d}$ Boolean functions. Therefore, the total number of AND gates is $2^{d}+2^{n-d}$, which has the claimed minimum value $O(\sqrt{2^{n}})$ by setting $d=n/2$. In the case of $b>1$, an analogous calculation shows the total number of AND gates is roughly $b\cdot 2^{d}+2^{n-d}$, which has minimum value $O(\sqrt{b\cdot 2^{n}})$ as claimed.
Note that the number of XOR gate used is $O(b\cdot 2^{n})$.
We shall also use the known T{T}-optimal synthesis of single-qubit unitaries without ancillas [27, 43, 28, 41].
@@
Any single-qubit unitary with determinant 11141414As noted in [43], this determinant condition is needed since a Clifford+T{T} circuit has determinant in $\{e^{i\cdot t\pi/4}\colon t=0,1,\ldots,7\}$. This is not a problem in practice because the determinant can be adjusted by multiplying the unitary with a global phase. can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using $O(\log(1/\varepsilon))$ T{T} gates and without ancillary qubits. [^15] In particular, the number of Clifford gates is also $O(\log(1/\varepsilon))$.
Below we give the proof of Theorem 1.2. The idea is to view the diagonal unitary $D$ as a single-qubit unitary acting on the $n$th qubit, controlled on the state of the first n1n-1 qubits. For each fixed value of the control bits, we have a single-qubit unitary that can be approximated by a sequence [^16] Since Lemma 2.3 uses no ancilla, the Clifford gates are only Hadamard H{H} and Phase gate $S=T^{2}$. Hence we simply assume the sequence contains only Hadamard and T{T} gates. of Hadamard and T{T} gates using Lemma 2.3. To implement $D$, we first use Lemma 2.1 to compute a Boolean function that takes as input the n1n-1 control bits, and outputs a description of the approximating sequence of Hadamard and T{T} gates into an ancilla register. Controlled on this ancilla register, we then apply the corresponding sequence of single-qubit gates.
@@
For each $j=1,2,\ldots,2^{n}$, define $U_{j}=\mathsf{diag}(\alpha_{j},\overline{\alpha}_{j})$, which is a single-qubit unitary with determinant $1$. Then $D^{\prime}=\mathsf{diag}(U_{1},\ldots,U_{2^{n}})$. Moreover, we know from Lemma 2.3 that each $U_{j}$ is $\varepsilon$-close to a single-qubit unitary $\widetilde{U}_{j}$ that can be exactly implemented as a sequence of $O(\log(1/\varepsilon))$ Hadamard and T{T} gates. It follows that $\|U-\widetilde{U}\|\leq\varepsilon$, where $\widetilde{U}=\mathsf{diag}(\widetilde{U}_{1},\ldots,\widetilde{U}_{2^{n}})$. To complete the proof we describe a Clifford+T{T} circuit that exactly implements $\widetilde{U}$ using the number of T{T} gates and ancillas claimed in the Theorem statement.
@@
Then by Lemma 2.1, the corresponding $(n+2K)$-qubit unitary $F$ mapping $|j\rangle|y\rangle$ to $|j\rangle|y\oplus f(j)\rangle$ can be implemented with $O\left(\sqrt{K\cdot 2^{n}}\right)=O\left(\sqrt{2^{n}\log(1/\varepsilon)}\right)$ T{T} gates and ancillas.
Once the gate sequence is computed into an ancilla register, it remains to apply H{H} or T{T} on the $(n+1)$-th qubit controlled on this ancilla register. For this purpose, we note that controlled-H{H} and controlled-T{T} are two-qubit unitaries that can be exactly implemented with $O(1)$ T{T} gates without ancillas [20].
@@
    For each $\ell=1,2,\ldots,2K$, we either apply H{H} on $\mathcal{A}_{1}$ controlled by $\mathcal{B}_{\ell}$ (if $\ell$ is odd), or we apply T{T} on $\mathcal{A}_{1}$ controlled by $\mathcal{B}_{\ell}$ (if $\ell$ is even).
@@
The number of T{T} gates used in the above implementation is upper bounded as
@@ 2.1 Batched synthesis
First we discuss batched synthesis of single-qubit unitaries: we will see that we can implement $m=O(\log\log(1/\varepsilon))$ single-qubit gates to error $\varepsilon$ with T{T}-count $O(\log(1/\varepsilon))$.
We are interested in synthesizing a general tensor product $U_{1}\otimes U_{2}\otimes\cdots\otimes U_{m}$ of single-qubit unitaries. In order to use Theorem 1.2 we will first reduce to the case where each $U_{j}$ is diagonal in the standard basis. To this end we can use the well-known Euler angle decomposition (see e.g., [36]) which expresses a single-qubit unitary $U$ as a product $U=e^{i\delta}e^{i\gamma{Z}}e^{i\beta{X}}e^{i\alpha{Z}}$ for some real numbers $\alpha,\beta,\gamma,\delta$. Here X,Y,Z{X},{Y},{Z} are the Pauli matrices. Since $e^{i\beta{X}}={H}e^{i\beta{Z}}{H}$ we can summarize this fact as follows.
@@
Let $U$ be a single-qubit unitary. Then there exist single-qubit diagonal unitaries $A,B,C$ such that U=AHBHCU=A{H}B{H}C, where H{H} is the single-qubit Hadamard matrix.
@@
Let $m\geq 1$ be an integer and $U_{1},U_{2},\ldots,U_{m}$ be arbitrary single-qubit unitaries. Define an $m$-qubit unitary $U=U_{1}\otimes\cdots\otimes U_{m}$. Then $U$ can be implemented up to error $\varepsilon$ by a Clifford+T{T} circuit using $O\left(\sqrt{2^{m}\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$ T{T} gates and ancilla qubits. [^17] This construction uses $O(2^{m}\log(1/\varepsilon))$ Clifford gates.
@@
By Fact 2.4, each $U_{j}$ can be expanded as $A_{j}{H}B_{j}{H}C_{j}$ where $A_{j},B_{j},C_{j}$ are single-qubit diagonal unitaries. As a result, we have $U=A{H}^{\otimes m}B{H}^{\otimes m}C$, where $A=A_{1}\otimes\cdots\otimes A_{m},B=B_{1}\otimes\cdots\otimes B_{m},C=C_{1}\otimes\cdots\otimes C_{m}$ are $m$-qubit diagonal unitaries, which can each be implemented using Theorem 1.2 with the claimed number of T{T} gates and ancillas. ∎
@@ 2.2 Mass production
Before proving this, let us describe our high-level strategy. We can use Fact 2.4 to reduce to the special case of implementing a diagonal single-qubit unitary $A:=\mathsf{diag}(1,e^{i\theta})$, by choosing the (irrelevant) global phase to be such that all 3 diagonal unitaries are of this form. For any computational basis input $|x\rangle$ where $x\in\{0,1\}^{m}$, we have $A^{\otimes m}|x\rangle=e^{i\theta\cdot|x|}|x\rangle$, where $|x|$ is the Hamming weight of $x$. A natural approach is then to first compute the Hamming weight into a second register containing $\lceil\log m\rceil$ qubits, then apply the corresponding phase controlled on this register. The Hamming weight can be computed using $O(m)$ gates, as we discuss below. In the second step we need to implement a diagonal unitary on $\lceil\log m\rceil$ qubits. We can do this using Theorem 1.2, incurring a T{T}-count $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)=O\left(m+\log(1/\varepsilon)\right)$.
@@
where $|x|\in\left\{0,1,\ldots,m-1\right\}$ is the Hamming weight of $x$. Then $U_{\mathsf{Ham}}$ can be exactly prepared by Clifford+T{T} circuit using $O(m)$ total gates and ancillary qubits.
We also remark that [11] gives the exact multiplicative complexity of computing the Hamming weight function, which upper bounds the T{T}-count of implementing $U_{\sf Ham}$ (using the connection in Remark 2.2). Their construction has slightly lower T-count than Fact 2.6.
@@
We describe a classical Boolean circuit with $O(m)$ gates to compute the Hamming weight of an $m$-bit string. Here the circuit contains AND, OR, NOT gates with bounded fan-in. Then we can convert this to a reversible classical circuit in the standard way using Toffoli gates and $O(m)$ ancilla qubits; implementing each Toffoli using $O(1)$ T{T} gates gives Fact 2.6.
Firstly note that given two $t$-bit numbers, we can compute their sum as a $(t+1)$-bit number using $O(t)$ gates, by the textbook algorithm sequentially computing each bit and the corresponding carry bit. The Boolean circuit to compute the Hamming weight will use the above observation iteratively for $t=1,2,\ldots,\log m$ as follows. We view the input as $m$ 11-bit numbers. Then we partition it into $m/2$ pairs and compute the sum of each pair as a $2$-bit number. Continuing in this way, at the $t$-th stage we will have $m/2^{t}$ $t$-bit numbers. The total gate complexity is
@@
By Fact 2.4, we obtain single-qubit diagonal unitaries $A,B,C$ such that U=AHBHCU=A{H}B{H}C. Thus $U^{\otimes m}=A^{\otimes m}{H}^{\otimes m}B^{\otimes m}{H}^{\otimes m}C^{\otimes m}$. Therefore it suffices to show how to implement $A^{\otimes m}$.
@@
    This uses $O(m)$ T{T} gates and ancillary qubits by Fact 2.6.
@@
    This uses $O\left(\sqrt{m\log(1/\varepsilon)}+\log(1/\varepsilon)\right)\leq O(m+\log(1/\varepsilon))$ T{T} gates and ancilla qubits by Theorem 1.2.
@@
    This again uses $O(m)$ T{T} gates and ancilla qubits by Fact 2.6.
Finally we prove the tightness of Theorem 1.4. On the one hand, the $\Omega(\log(1/\varepsilon))$ T{T}-count lower bound holds even if $m=1$ [5]. On the other hand, consider the special case where $U$ is the T{T} gate. Then $U^{\otimes m}={T}^{\otimes m}$ and it is impossible to implement it using $o(m)$ T{T} gates. Otherwise, by gate injection $|{T}\rangle^{\otimes m}$ can be approximated with T{T}-count $o(m)$. Plugging this in Theorem 1.1, we have that any $n$-qubit state can be approximated up to any constant error with T{T}-count $o(\sqrt{2^{n}})$, contradicting the lower bound for Theorem 1.1.181818The lower bound for Theorem 1.1 holds in the presence of Pauli postselections, as needed here after gate injection. See details in Section 4.